5. Smoking and duodenal ulcers. The Health Professionals Follow-up Study is a prospective study of heart disease and cancer among more than 50,000 health professionals in the United States who were 40–75 years of age in 1986. Every two years questionnaires are sent to these individuals, and newly diagnosed cases of various diseases reported. The following data are constructed from the surveys returned in the 1993 mailing. The investigators in this study were interested in the relationship between smoking status and duodenal ulcers, a common disorder of the gastrointestinal tract. The incidence of duodenal ulcers for the three groups is presented below:

Smoking and the Incidence of Duodenal Ulcers

Number of Persons at Start of study

Number of Observed New Cases of Duodenal Ulcers

Nonsmokers

22,295

60

Past smokers

20,757

60

Current smokers

4,754

16

Calculate the relative risk of being a past smoker and a current smoker, relative to never having smoked. (Hint: You should make two tables, one for smokers and one for past smokers.)
Answer #1
The relative risk between current smokers and nonsmokers
Number of Persons at Start of study Number of Observed New Cases of Duodenal Ulcers Total
Current smokers 4,754 16 4770
Nonsmokers 22,295 60 22,355
Total 27,049 76

The incidence of exposed is 4754: (4754 +16) = 0.996

The Incidence of non exposed 22. 295 : (22, 295 +60) = 0.997

So, the Relative Risk among current smokers and nonsmokers is

0.996 0.997 = 0,998

The risk is almost same for both group

its means there is no significant chance of duodenal ulcer among current smokers than none smokers.

There is no association between current smoking and duodenal ulcer

The relative risk between past smokers and nonsmokers
Number of Persons at Start of study Number of Observed New Cases of Duodenal Ulcers Total
Past smokers 20,757 60 20,817
Nonsmokers 22,295 60 22.355
Total 43,052 120

The incidence of exposed is 20.757 : (20.757 +60) = 0.997

The incidence of non exposed is 22. 295 : (22, 295 +60) = 0.997

So, the Relative Risk among past smokers and nonsmokers is

0.997:0.997=1

the risk is the same for both of group

So, there is no association between Past smoking and duodenal ulcer